A basketball player scored a total of 19 points for making 7 baskets. He made 3 and 2 points only. How many of each type of of shots did he make?
Answer by JulietG(1812)

If the player made all 3-point shots, his total would be 7*3, or 21
His actual total is 19.
The difference is 2.
The difference between a 3 point shot and a 2 point shot is 1 point.
2/1 = 2 two-point shots
That leaves 7-2, or 5 three-pointers
(5) 3-pointers = 15
(2) 2-pointers = 4
15+4=19
Success!